# Transformers for Stratified Spectropolarimetric Inversion: Proof of Concept

### v2

@@ Appendix A Azimuth predictions
Figures 15 and  16 show the azimuthal predictions for the MLP and transformer. In most of the snapshot, for both ML models, the azimuth has relatively large errors. The darkest regions in the residual maps do indicate however that there are distinct regions where the azimuth has been successfully predicted, and consistent with the other atmospheric parameters, the transformer clearly performs more effectively at this task than the MLP. In order for the azimuth to be constrained (without disambiguation), there must be sufficiently large polarisation amplitudes in Stokes $Q$ and $U$, which occurs in only a small minority of pixels due to the quiet Sun nature of the MANCHA snapshots.

### v3

@@ Appendix A Azimuth predictions
Figures 17 and  18 show the azimuthal predictions for the MLP and transformer. In most of the snapshot, for both ML models, the azimuth has relatively large errors. The darkest regions in the residual maps do indicate however that there are distinct regions where the azimuth has been successfully predicted, and consistent with the other atmospheric parameters, the transformer clearly performs more effectively at this task than the MLP. In order for the azimuth to be constrained (without disambiguation), there must be sufficiently large polarisation amplitudes in Stokes $Q$ and $U$, which occurs in only a small minority of pixels due to the quiet Sun nature of the MANCHA snapshots.